Need help solving this:

Complex analysis.

\int_{0}^{\infty}\frac{sin(2x)}{x(x^2+4)}dx
Answer #1

1 Solution het fiza iaz 2(22+4) Consider the Pintegran ffres da coleu c cn the closed . contoor consisting of the upper of th2=0, Izi fiz, has only one bole of ander 1 at z = ai cwithin The pole z=0 in aroided by identation By cechys Gresidues theor13 eidz clicy z fiz? = lim zo (22+4) Zo Therefore, liu fezidz = -i (T-0) 4 AB Ho 8 ill an are & Cosp of the carde Izlar liveidz Now fazla مهام ZlZ²+4) Z(2-21)(< tai) iaz e - Res(222i) Lt (2-21). Zaai Z12-21) (2+2 i) lalai) clar e = = Lt 2+2i Z(2+21)Equeeting imaginary parts, are 6 how sas Sinax dx= TT u Tie x(2²+4) 4 Sinan x(x²+4)